# Rejection-Sampled Linear Codes for Lossy Compression and Channel Simulation

### v1

@@ VIII Mixed GRSSE and Experiments
The $\mathbf{H}^{(1)}/\cdots/\mathbf{H}^{(N)}$-GRSSE, or the mixed GRSSE, is the GRSSE where each $\mathbf{H}_{i}$ is selected within a prescribed sequence of codes $\mathbf{H}^{(1)},\ldots,\mathbf{H}^{(N)}$ in this order, i.e., $(\mathbf{H}_{i})_{i}$ is formed by repeating $\mathbf{H}^{(1)}$ a nonnegative integer number of times, then $\mathbf{H}^{(2)}$ a nonnegative integer number of times, etc. The precise number of times can be determined by a computer search that attempts to minimize the expected communication cost. A simple efficient rule that performs well in experiments is to simply select the code $\mathbf{H}^{(j_{i})}$ at iteration $i$ via
@@
where $\epsilon>0$ is a small fixed parameter, and $\tilde{d}_{j}$ is the distance of the code $\mathbf{H}^{(j)}$ (or an estimate of the distance), assuming that $\tilde{d}_{1}>\cdots>\tilde{d}_{N}$. The reason is that Theorem 6 suggests that a smaller $\mathbb{P}(\mathrm{wt}(\mathbf{Z})\geq d/2)$ would result in a smaller communication cost. This rule shares some similarities with the adaptive greedy rejection sampling scheme [9], which also adaptively changes the proposal distribution at each iteration.
We discuss another reason to consider mixed GRSSE. A limitation of the $\mathbf{H}$-GRSSE is that the precomputation of $\kappa_{i}$ in Algorithm 2 may not terminate, i.e., the number of iterations is unbounded. Even though Algorithm 2 does not terminate, the actual encoding and decoding in Algorithm 1 almost surely terminates since $\mathbb{E}[\log L]<\infty$ by Theorem 6. In practice, we can run Algorithm 2 in a “lazy” manner, where we only compute $\kappa_{i}$ when the encoding algorithm in Algorithm 1 reaches iteration $i$. This way, we can ensure a finite running time for $\mathbf{H}$-GRSSE. Nevertheless, if we require the precomputation in Algorithm 2 to terminate in finite time (e.g., if we want to compute the expected communication cost precisely), we can consider the $\mathbf{H}/\text{complete}$-GRSSE, where $\mathbf{H}^{(1)}=\mathbf{H}$ and $\mathbf{H}^{(2)}$ is the empty matrix (the complete code). When the encoder reachs the iteration where $\mathbf{H}_{i}$ is the complete code, it will simply send $\mathbf{M}=\mathbf{X}$ and terminate the algorithm.
@@
    The mixed-GRSSE algorithm using the rule (7), with $\epsilon=10^{-9}$ and $\mathbf{H}^{(1)},\ldots,\mathbf{H}^{(N)}$ being the best known $[24,k]$ linear block codes [47] with the largest distance for $k=0,1,2,3,5,7,12,14,18,19,23,24$, computed using Magma [48].
@@
Since the Golay code has a distance $8$, according to Theorem 6, Golay-GRSSE is only suitable when $\mathbb{P}(\mathrm{wt}(\mathbf{Z})\leq 3)$ is large, meaning that it requires $\alpha\leq 3/24$. Practically, $\alpha$ has to be lower to ensure $\mathrm{wt}(\mathbf{Z})\leq 3$ most of the time. We can see from Figure 3 that Golay-GRSSE is suitable for $\alpha\leq 0.1$, where its rate is close to the excess functional information lower bound. In comparison, the mixed-GRSSE is close to the lower bound for every $\alpha$, due to the ability to choose a code with a suitable distance at each iteration. The gap between the rate of mixed-GRSSE and the excess functional information lower bound is at most $0.06$ bits.
@@
IX GRSSE with Juxtaposition Codes
We now discuss a method to apply GRSSE with a larger blocklength $n$. A simple way to increase the blocklength of a code is to perform juxtaposition, i.e., a codeword in the resultant code is the juxtaposition of any $r$ codewords of the original code. The parity-check matrix of the juxtaposition code is the block-diagonal matrix $\mathbf{I}_{r}\otimes\mathbf{H}$, where where $\mathbf{H}$ is the parity-check matrix of the original code, $\mathbf{I}_{r}$ is the $r\times r$ identity matrix, and $\otimes$ is the Kronecker product. The type set distribution $p_{t_{\mathbf{I}_{r}\otimes\mathbf{H}}}$ (3) of the juxtaposition code $\mathbf{I}_{r}\otimes\mathbf{H}$ can be computed efficiently in terms of the type set distribution $p_{t_{\mathbf{H}}}$ of the original code, by noting that $p_{t_{\mathbf{I}_{r}\otimes\mathbf{H}}}$ is the distribution of $\{r^{-1}\sum_{i=1}^{r}\tau_{i}:\tau_{1}\in t_{1},\ldots,\tau_{r}\in t_{r}\}$ where $t_{i}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}p_{t_{\mathbf{H}}}$, i.e., $p_{t_{\mathbf{I}_{r}\otimes\mathbf{H}}}$ is the convolution (with respect to Minkowski sum) of $r$ copies of $p_{t_{\mathbf{H}}}$.
@@ IX GRSSE with Juxtaposition Codes
    The 4×4$\times$4 ×Golay-GRSSE with $n=96$, which is the pure GRSSE with inner code $\mathbf{I}_{4}\otimes\mathbf{H}$ being the the juxtaposition of four $[24,12,8]$ Golay codes. [^4] For the sake of efficiency, the number of iterations for both $\mathbf{I}_{4}\otimes\mathbf{H}$-GRSSE and mixed-GRSSE is capped at $20000$, i.e., we switch to the complete code at iteration $20000$.
@@
    The mixed-GRSSE with $n=96$ using the rule (7), with $\epsilon=0.01$, $\mathbf{H}^{(1)},\ldots,\mathbf{H}^{(N)}$ being the trivial code ($k=0$), the repetition code ($k=1$), and the $4$-fold juxtaposition of the best known $[24,k]$ linear block codes [47] with the largest distance for $k=12,14,18,19,23,24$, computed using Magma [48]. [^5] We choose the better one among this rule and fixing the code to be the $4$-fold juxtaposition of the Golay code (same as the 4×4$\times$4 ×Golay-GRSSE). For the $r$-fold juxtaposition of a code with distance $d_{j}$, we use an estimated distance $\tilde{d}_{j}=rd_{j}$ in rule (7). [^6] Technically, the distance of the $r$-fold juxtaposition is still $d_{j}$. Nevertheless, there are only $r$ codewords with weight $d_{j}$, and their contribution is small in the GRSSE algorithm, which degrades gracefully when a small number of codewords have weights that are too small. Intuitively, the “effective distance” of the $r$-fold juxtaposition is $\tilde{d}_{j}=rd_{j}$, and this estimate performs well in the experiment.
@@
Figure 5 shows that 4×4$\times$4 ×Golay-GRSSE performs well when $\alpha\in[0.1,0.125]$. We can also see that mixed-GRSSE performs well when $\alpha\leq 0.125$. PolarSim achieves a smaller communication rate, but requires a significantly larger $n$ to have a satisfactory performance. The difference between the rates of mixed-GRSSE ($n=96$) and PolarSim ($n=4096$) is small when $\alpha\leq 0.125$ (the difference is at most $0.058$ bits).

### v2

@@ IX Mixed GRSSE and Experiments
The $\mathbf{H}^{[1]}/\cdots/\mathbf{H}^{[N]}$-GRSSE, or the mixed GRSSE, is the GRSSE where each $\mathbf{H}_{i}$ is selected within a prescribed sequence of codes $\mathbf{H}^{[1]},\ldots,\mathbf{H}^{[N]}$ in this order, i.e., $(\mathbf{H}_{i})_{i}$ is formed by repeating $\mathbf{H}^{[1]}$ a nonnegative integer number of times, then $\mathbf{H}^{[2]}$ a nonnegative integer number of times, etc. The precise number of times can be determined by a computer search that attempts to minimize the expected communication cost. A simple efficient rule that performs well in experiments is to simply select the code $\mathbf{H}^{[j_{i}]}$ at iteration $i$ via
@@
where $\epsilon>0$ is a small fixed parameter, and $\tilde{d}_{j}$ is the distance of the code $\mathbf{H}^{[j]}$ (or an estimate of the distance), assuming that $\tilde{d}_{1}>\cdots>\tilde{d}_{N}$. The reason is that Theorem 15 suggests that a smaller $\mathbb{P}(\mathrm{wt}(\mathbf{Z})\geq d/2)$ would result in a smaller communication cost. This rule shares some similarities with the adaptive greedy rejection sampling scheme [9], which also adaptively changes the proposal distribution at each iteration.
We discuss another reason to consider mixed GRSSE. A limitation of the $\mathbf{H}$-GRSSE is that the precomputation of $\kappa_{i}$ in Algorithm 2 may not terminate, i.e., the number of iterations is unbounded. Even though Algorithm 2 does not terminate, the actual encoding and decoding in Algorithm 1 almost surely terminates since $\mathbb{E}[\log L]<\infty$ by Theorem 15. In practice, we can run Algorithm 2 in a “lazy” manner, where we only compute $\kappa_{i}$ when the encoding algorithm in Algorithm 1 reaches iteration $i$. This way, we can ensure a finite running time for $\mathbf{H}$-GRSSE. Nevertheless, if we require the precomputation in Algorithm 2 to terminate in finite time (e.g., if we want to compute the expected communication cost precisely), we can consider the $\mathbf{H}/\text{complete}$-GRSSE, where $\mathbf{H}^{[1]}=\mathbf{H}$ and $\mathbf{H}^{[2]}$ is the empty matrix (the complete code). When the encoder reachs the iteration where $\mathbf{H}_{i}$ is the complete code, it will simply send $\mathbf{M}=\mathbf{X}$ and terminate the algorithm.
@@
    The mixed-GRSSE algorithm using the rule (14), with $\epsilon=10^{-9}$ and $\mathbf{H}^{[1]},\ldots,\mathbf{H}^{[N]}$ being the best known $[24,k]$ linear block codes [56] with the largest distance for $k=0,1,2,3,5,7,12,14,18,19,23,24$, computed using Magma [57].
@@
Since the Golay code has a distance $8$, according to Theorem 15, Golay-GRSSE is only suitable when $\mathbb{P}(\mathrm{wt}(\mathbf{Z})\leq 3)$ is large, meaning that it requires $\alpha\leq 3/24$. Practically, $\alpha$ has to be lower to ensure $\mathrm{wt}(\mathbf{Z})\leq 3$ most of the time. We can see from Figure 4 that Golay-GRSSE is suitable for $\alpha\leq 0.1$, where its rate is close to the excess functional information lower bound. In comparison, the mixed-GRSSE is close to the lower bound for every $\alpha$, due to the ability to choose a code with a suitable distance at each iteration. The gap between the rate of mixed-GRSSE and the excess functional information lower bound is at most $0.06$.
@@
X GRSSE with Juxtaposition Codes
We now discuss a method to apply GRSSE with a larger blocklength $n$. A simple way to increase the blocklength of a code is to perform juxtaposition, i.e., a codeword in the resultant code is the juxtaposition of any $r$ codewords of the original code. The parity-check matrix of the juxtaposition code is the block-diagonal matrix $\mathbf{I}_{r}\otimes\mathbf{H}$, where where $\mathbf{H}$ is the parity-check matrix of the original code, $\mathbf{I}_{r}$ is the $r\times r$ identity matrix, and $\otimes$ is the Kronecker product. The type set distribution $p_{\mathcal{T}_{\mathbf{I}_{r}\otimes\mathbf{H}}}$ (10) of the juxtaposition code $\mathbf{I}_{r}\otimes\mathbf{H}$ can be computed efficiently in terms of the type set distribution $p_{\mathcal{T}_{\mathbf{H}}}$ of the original code, by noting that $p_{\mathcal{T}_{\mathbf{I}_{r}\otimes\mathbf{H}}}$ is the distribution of $\{r^{-1}\sum_{i=1}^{r}\mathbf{t}_{i}:\mathbf{t}_{1}\in\mathcal{T}_{1},\ldots,\mathbf{t}_{r}\in\mathcal{T}_{r}\}$ where $\mathcal{T}_{i}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}p_{\mathcal{T}_{\mathbf{H}}}$, i.e., $p_{\mathcal{T}_{\mathbf{I}_{r}\otimes\mathbf{H}}}$ is the convolution (with respect to Minkowski sum) of $r$ copies of $p_{\mathcal{T}_{\mathbf{H}}}$.
@@ X GRSSE with Juxtaposition Codes
    The $4\timesGolay$-GRSSE with $n=96$, which is the pure GRSSE with inner code $\mathbf{I}_{4}\otimes\mathbf{H}$ being the the juxtaposition of four $[24,12,8]$ Golay codes. [^5] For the sake of efficiency, the number of iterations for both $\mathbf{I}_{4}\otimes\mathbf{H}$-GRSSE and mixed-GRSSE is capped at $20000$, i.e., we switch to the complete code at iteration $20000$.
@@
    The mixed-GRSSE with $n=96$ using the rule (14), with $\epsilon=0.01$, $\mathbf{H}^{(1)},\ldots,\mathbf{H}^{(N)}$ being the trivial code ($k=0$), the repetition code ($k=1$), and the $4$-fold juxtaposition of the best known $[24,k]$ linear block codes [56] with the largest distance for $k=12,14,18,19,23,24$, computed using Magma [57]. [^6] We choose the better one among this rule and fixing the code to be the $4$-fold juxtaposition of the Golay code (same as the $4\timesGolay$-GRSSE). For the $r$-fold juxtaposition of a code with distance $d_{j}$, we use an estimated distance $\tilde{d}_{j}=rd_{j}$ in rule (14). [^7] Technically, the distance of the $r$-fold juxtaposition is still $d_{j}$. Nevertheless, there are only $r$ codewords with weight $d_{j}$, and their contribution is small in the GRSSE algorithm, which degrades gracefully when a small number of codewords have weights that are too small. Intuitively, the “effective distance” of the $r$-fold juxtaposition is $\tilde{d}_{j}=rd_{j}$, and this estimate performs well in the experiment.
@@
Figure 6 shows that $4\timesGolay$-GRSSE performs well when $\alpha\in[0.1,0.125]$. We can also see that mixed-GRSSE performs well when $\alpha\leq 0.125$. PolarSim achieves a smaller communication rate, but requires a significantly larger $n$ to have a satisfactory performance. The difference between the rates of mixed-GRSSE ($n=96$) and PolarSim ($n=4096$) is small when $\alpha\leq 0.125$ (the difference is at most $0.058$).